# Logic-in-Frames: Dynamic Keyframe Search via Visual Semantic-Logical Verification for Long Video Understanding

### v1

@@ Abstract
Figure 1: Examples of four types of visual semantic-logical relationships in video QA detected by our VSLS: Time (text, time, pen), Attribute (man, attribute, white shirt), Spatial (copilot, spatial, Egyptian Pyramids), and Causal (man, causal, basketball). Green boxes indicate correct answers, while red boxes show baseline errors.
@@ 1 Introduction
Figure 2: Visual Semantic-Logical Search (VSLS) Framework for Efficient Keyframe Selection in Long Video Understanding. The figure illustrates our VSLS framework for efficient keyframe search in long video understanding tasks. The pipeline begins with sparse sampling from the input video, followed by iterative frame selection based on object detection and semantic logical relation verification. Key components include: 1) Query decomposition via LLM/VLM to extract cue objects, target objects, and four types of logical relations (spatial, temporal, attribute, and causal); 2) Adaptive grid sampling that selects frames based on an evolving confidence distribution; 3) Object detection using YOLO-World to identify query-relevant entities; 4) Multi-faceted confidence estimation combining object and logic relation confidence scores through a spline function; selected high-confidence frames are fed to downstream tasks.
@@
Vision-Language Models (VLMs) [40] have achieved remarkable progress in visual understanding [47, 29], particularly in video understanding [32, 44], which holds significant potential for modeling real-world scenarios. However, effectively modeling raw video data remains challenging due to the high-dimensional nature of spatio-temporal information and inherent model complexity constraints. These limitations severely restrict temporal dependency modeling, confining existing approaches to processing only a limited number of frames simultaneously - a critical bottleneck for comprehensive video understanding. This paper addresses a pivotal research question: How can we efficiently and accurately select keyframes that are semantically critical for answering video-based queries?
@@
We hypothesize that deconstructing visual semantic and logical cues (e.g., target objects, logical relations between visual entities) from textual queries enables effective identification of task-relevant frames through heuristic sampling and search. Building on this insight, we propose Visual Semantic-Logical Search (VSLS), a novel framework that iteratively selects minimal yet maximally informative frames through target object confidence estimation and joint verification of visual semantic logic. Our method achieves temporal localization through iterative optimization of the sampling distribution. This approach requires only sparse sampling (1.4%) to identify critical frames while maintaining the temporal and spatial coherence of visual entities, significantly reducing computational complexity compared to conventional dense sampling strategies.
Compared with existing methods, VSLS offers three unique advantages. First, VSLS enables highly efficient frame retrieval without dense captioning (e.g., [4, 15, 31]) or video clustering (e.g., [33, 24]). By iteratively refining a sampling distribution, VSLS achieves precise temporal localization with only sampling 1.4% frames on average of LVHaystack. Second, VSLS comprehensively accounts for logical relations in the query, going beyond simple target detection (e.g., [39]). Specifically, it supports four types of relational logic—spatial co-occurrence, temporal dependencies, attribute dependencies, and causal relations—allowing it to accommodate a broader range of query types. Third, VSLS functions as a plug-and-play module independent of both the query analysis LLM and the QA-oriented VLM, enabling integration into existing pipelines without interdependency among components.
@@
We examined VSLS on several public datasets, including LongVideoBench [38], a comprehensive benchmark for evaluating models on long-form video understanding tasks; Video-MME [11], a widely adopted multimodal evaluation benchmark for video understanding; and Haystack-LVBench [38], which features meticulously annotated frame-level labels for precise temporal analysis. Our extensive experiments demonstrate significant improvements in the semantic similarity between our retrieved keyframes and ground truth, temporal coverage, and accuracy on downstream long video QA tasks. More importantly, by sampling only 1.4% of video frames (Ego4D [12]) for detection and updating the sampling distribution, we achieved an 8.7% improvement in GPT-4o [14]’s long video QA accuracy. This result is attributed to our simple yet powerful observation: query-guided visual semantic logic retrieval can mitigate the gap between potential visual logic and the logic expressed in the query text. These findings suggest that constructing ternary logic triplets with visual elements (e.g., object1, logic type, object2) can enhance downstream reasoning capabilities when performing text-visual retrieval.
@@
To our best knowledge, we are arguably the first to search for keyframes in long videos by detecting visual semantic logic, with potential extensions to other text-vision retrieval tasks. To summarize, we make the following contributions:
@@
  • We define four prevalent types of semantic logic relations in video QA tasks, including temporal, causal, attribute, and spatial relations, which can be accurately detected.
@@
  • We require sampling only 1.4% of frames on average to complete the keyframe search through (1) heuristic sampling and (2) updating the sampling distribution according to different video semantics and logical relation.
@@
  • We validate several metrics, including retrieval efficiency, semantic similarity, and temporal coverage on multiple widely used open-source datasets, demonstrating significant improvements in downstream tasks.
@@ 2 Method
Although existing long-context VLM frameworks implement keyframe search for video QA tasks [17, 21, 28, 31, 33, 41], their computational efficiency and searching accuracy remain suboptimal. To address this needle-in-a-haystack challenge [30, 46], we propose a novel method VSLS that aligns the semantic relations between the text modality and video modality, enhancing the plausibility of logical reasoning and performance of downstream tasks.
@@ 2.1 Task Formulation
Given a video sequence $V=\{f_{t}\}_{t=1}^{N_{v}}$ with $N_{v}$ frames and query $Q$, the temporal search aims to determine the minimal keyframe subset $V^{K}=\{f_{m_{i}}\}_{i=1}^{m}\subseteq V$ that satisfies:
@@
  • Compactness: $V^{K}$ must be a minimal subset preserving completeness, meaning no frame can be removed without affecting answer consistency, also ensuring efficiency.
@@ 2.2 Visual Semantic Logic Extraction
Given a question $Q$ and uniform sampled frames $\overline{V}_{N}$ from a video $V$, our goal is to identify and organize the essential visual elements for answering $Q$. To this end, we categorize detected items into two groups:
@@
To reason about semantic and logical links among these objects, we define a set of relations $\mathcal{R}\subseteq\mathcal{O}\times\Delta\times\mathcal{O}$, where each relation $r\in\mathcal{R}$ is represented as an ordered triple $(o_{i},\delta,o_{j})$, with $\delta\in\Delta$ denoting one of the following types:
  • Spatial co-occurrence: $o_{i}$ and $o_{j}$ must appear in the same frame, reflecting co-occurrence or proximity, e.g., for sentence “A person is standing beside a vase.” the relation will be (person, spatial, vase).
  • Attribute Dependency: $o_{i}$ and $o_{j}$ share a descriptive property such as color or size, capturing appearance-based attributes, e.g., for sentence “A person wears a black shirt.” the relation will be (person, attribute, black shirt).
  • Temporal Proximity: $o_{i}$ and $o_{j}$ occur in different frames within a short time interval, linking event progressions or changing states, e.g., for sentence “After a dog entered the room, a cat entered the room.” the relation will be (dog, temporal, cat).
  • Causal Order: $o_{i}$ and $o_{j}$ follow a clear temporal order indicating a cause–effect or prerequisite relation, e.g., for sentence “A little girl broke the vase.” the relation will be (little girl, causal, pieces).
These four types are chosen to comprehensively capture spatial co-occurrence, attribute dependency, temporal proximity, and cause-effect structures. As shown in Figure 1, we form the semantic-logical relation that effectively addresses a wide range of question-answering tasks. In particular, questions requiring temporal queries (“when does X happen?”), causal reasoning (“why did Y occur?”), attribute dependence (“What is the person wearing sunglasses doing?”) or detailed spatial constraints (“Who is standing next to the red car?”) can be more reliably answered by integrating these structured relations and contextual cues.
@@ 2.3 Iterative Semantic-Logical Temporal Search
Our algorithm iteratively searches for keyframes by semantic and logical reasoning, consisting of four main components: Frame Sampling, Object Detection and Scoring, Visual Semantic Logic Detection, and Distribution Update. The full pseudocode is provided in Algorithm 1.
1
2Function SemanticLogicalTemporalSearch($V,Q,K,\Delta_{t},\tau,\alpha,\gamma$)
3       $\mathcal{O},\mathcal{R}\leftarrow\text{ParseQuestion}(Q)$
4       // Extract key/cue objects and relations
5      $P\leftarrow\text{Uniform},B\leftarrow|V|,S\leftarrow\emptyset,N_{v}\leftarrow|V|$
6       // Initialize distribution and state
7      while $B>0$ and $|\mathcal{O}|>0$ do
8             $k\leftarrow\lfloor\sqrt{B}\rfloor$, $G\leftarrow\textnormal{{Grid}}(\textnormal{{Sample}}(P,k^{2}))$
9             // Adaptive grid sampling
10            $\Omega\leftarrow\textnormal{{DetectObjects}}(G)$
11             // Detect objects in sampled frames
12            foreach $t\in G$ do
13                   $C_{t}\leftarrow\textnormal{{CalculateBaseScore}}(\Omega[t])$
14                   // Base detection confidence
15                  foreach  $r_{type}\in\mathcal{R}$  do
16                         $\delta\leftarrow\textnormal{{Processrelation}}(r_{type},\Omega,\Delta_{t},\tau,\alpha,\gamma)$
17                         // relations require distinct processing
18                        $C_{t}\leftarrow C_{t}+\delta$
19                  $\textnormal{{UpdateScores}}(S,t,C_{t})$
20                   // Update global score registry
21
22            $\textnormal{{DiffuseScores}}(S,w)$
23             // Temporal context propagation
24            $P\leftarrow\textnormal{{NormalizeDistribution}}(S)$, $B\leftarrow B-k^{2}$
25             // Update sampling distribution
26            foreach $g\in\textnormal{{TopK}}(S,K)$ do
27                   if $\Omega[g]\cap\mathcal{O}\neq\emptyset$ then
28                         $\mathcal{O}\leftarrow\mathcal{O}\setminus\Omega[g]$
29                         // Remove identified key objects
30
31
32      return $\textnormal{{TopK}}(S,K)$
33       // Return top-K keyframes
@@ 2.3.1 Frame Sampling
To accelerate the search process, we do not examine each frame of the video exhaustively; instead, we adopt a distributed sampling strategy. Let $N_{v}$ denote the total number of frames in the video, and let $P$ be a sampling distribution initialized to uniform over all frames. We define the sampling process as:
@@
where $\mathrm{Sample}(\cdot,N_{s})$ selects a subset of $N_{s}$ frames according to the distribution $P\odot N_{v}$. To further accelerate subsequent steps, we stack the sampled frames into a $k\times k$ grid, which imposes a constraint on the sample size $N_{s}$. Specifically, we require:
@@ 2.3.2 Object Detection and Scoring
Method Training Required Searching Efficiency Overall Task Efficiency Matching Iteration TFLOPs $\downarrow$ Latency (sec) $\downarrow$ Latency (sec) $\downarrow$ Acc $\uparrow$ Static Frame Sampling Uniform-8 [38] Training-Based N/A N/A N/A 0.2 3.8 53.7 Dense Retrieval VideoAgent [10] Training-Based CLIP-1B [23] 840 536.5 30.2 34.9 49.2 T*-Retrieval [39] Training-Based YOLO-World-110M 840 216.1 28.6 32.2 57.3 Temporal Search T*-Attention [39] Training-Based N/A N/A 88.9 13.7 17.3 59.3 T*-Detector [39] Training-Free YOLO-World-110M 43 31.7 7.3 11.1 59.8 VSLS (ours)-Detector Training-Free YOLO-World-110M 49 33.3 7.8 11.6 61.5
Table 1: Evaluation of computational performance metrics across the LV-Haystack benchmark, presenting both search efficiency and end-to-end processing overhead (combining search and inference stages).
In this stage, we construct the detection search space by taking the union of both Key Objects and Cue Objects. For each iteration, we detect objects on the $N_{s}$ sampled frames using a lightweight model like YOLO-World [6] for high efficiency and score the frames based on detection confidence. Specifically, let $D_{t}$ be the set of detected objects in the frame at time $t$, $c_{o}$ the confidence of each detected object, and $w_{o}$ the corresponding weight. We define the frame score as:
$C_{t}=\max_{o\in D_{t}}\bigl{(}c_{o}\cdot w_{o}\bigr{)}.$ (3)
@@ 2.3.3 Visual Semantic Logic Detection
Beyond individual object detection and frame-level scoring, we further refine each frame’s confidence score by modeling higher-order logic relations among objects. Let $\mathcal{R}$ be a set of relations, where each relation $r\in\mathcal{R}$ is defined on a pair of objects $(o_{1},o_{2})$ and labeled with a relation type $r_{\text{type}}$. Denote by $C_{t}$ the confidence score for the frame at time $t$, and let $\alpha$ be a global scaling factor shared by all relations. We introduce an additional relation-specific weight $\gamma_{r_{\text{type}}}$ to control how strongly each type of logic affects the frame score. Formally, the updated frame confidence $C_{t}^{(r)}$ after satisfying relation $r$ is given by:
@@
Attribute Relation. An attribute relation is satisfied when $o_{1}$ and $o_{2}$ share sufficient bounding-box overlap in the same frame. Let $\mathrm{overlap}$ be the ratio of their intersection area to the minimum of their individual bounding-box areas. If $\mathrm{overlap}>\tau$, we increment the frame confidence:
@@
Time Relation. A time relation checks whether two objects appear in temporally close frames. Suppose $t_{i}$ and $t_{j}$ ($t_{i}\leq t_{j}$) are sampled such that $|\,t_{j}-t_{i}|<\Delta_{t}$, where $\Delta_{t}$ is a threshold (e.g. five frames), if $o_{1}$ occurs in frame $t_{i}$ and $o_{2}$ in frame $t_{j}$, then both frames’ confidences are updated:
@@
Through these steps, frames that fulfill any relation criteria receive an incremental boost in confidence, thereby emphasizing contextually consistent detections. As we accumulate evidence from spatial, temporal, causal, and attribute relations, the system more accurately identifies the keyframes and interactions in the video.
2.4 Distribution Update
After each iteration of frame sampling, we merge newly obtained frame confidences into the global score distribution $\{S_{f}\}$ spanning all frames $f=1,2,\ldots,N_{v}$. When a frame $f$ is selected for detection, its score is assigned to the confidence value $C_{f}$, and the visitation counter $N_{v,f}$ is reset:
[…]
To incorporate temporal context, we diffuse this updated score to neighboring frames within a window of size $w$. Denoting each nearby index by $f\pm\delta$ (for $\delta\in[-w,w]$), we apply:

### v2

@@ 1 Introduction
Vision-Language Models (VLMs) yin2024survey have achieved remarkable progress in video understanding zou2024seconds; tang2023video, particularly in video question answering wang2024gpt4video; 84, demonstrating potential for modeling real-world scenarios. However, existing methods can only simultaneously process a limited number of frames due to the inherent token limit and extremely high dimension of spatio-temporal video data, especially for long videos. Furthermore, uniformly sampled keyframes are query-agnostic and insufficient to represent query-related contents. To tackle these challenges, this paper addresses a pivotal research question:
How can we efficiently and accurately select keyframes that are semantically critical for answering video-based queries?
@@
We hypothesize that deconstructing visual semantic and logical cues (e.g., target objects, logical relations including temporal, spatial, attribute, and causal relationships between visual entities) from textual queries enables effective identification of task-relevant frames through heuristic sampling and search. Building on this insight, we propose Visual Semantic-Logical Search (VSLS), a novel keyframe search framework that incorporates target object confidence estimation and joint verification of visual semantic logic into the iterative update of frame sampling distribution and selects the most informative frames with the highest confidence. Experimental results show that our approach requires only sparse sampling (1.4% of frames per video on average) to identify critical frames, significantly reducing computational complexity compared to conventional dense sampling strategies while maintaining performance on downstream video understanding tasks.
Figure 1: Examples of four types of visual semantic-logical relationships in video QA detected by our VSLS framework: Temporal (text, time, pen), Attribute (man, attribute, white shirt), Spatial (copilot, spatial, Egyptian Pyramids), and Causal (man, causal, basketball). Green boxes indicate correct answers, while red boxes show baseline errors.
Compared to conventional methods, VSLS shows three distinct advantages. First, the framework is training-free and highly efficient in comparison with dense captioning chen2024sharegpt4video; kim2024you; wang2024videoagent or video clustering wang2024videotree; rajan2025key strategies, sampling only 1.4% of frames on average in LVHaystack. Second, it explicitly models logical binary relations (namely spatial, temporal, attribute, and causal) in the query beyond simple target detection ye2025rethinking, utilizing additional visual semantic features and enhancing logical consistency throughout the reasoning process. Third, VSLS is a plug-and-play module, which can be seamlessly integrated into existing VLM pipelines without cross-component dependencies.
@@
We further examine VSLS on several public datasets, including LongVideoBench lvhdataset, a comprehensive benchmark for long video understanding; Video-MME videomme, a widely adopted multimodal video question answering dataset; and Haystack-LVBench lvhdataset with meticulously annotated keyframes based on human feedback for more precise analysis. Extensive experiments demonstrate significant improvements in both the semantic similarity and temporal coverage between the retrieved keyframes and the ground truth labels, as well as the accuracy in downstream video question-answering tasks. More importantly, with only 1.4% of video frames (Ego4D grauman2022ego4d) sampled in the search iteration, our method achieves an 8.7% improvement in GPT-4o hurst2024gpt’s long video QA accuracy. This performance gain is attributed to our simple yet powerful observation: query-guided visual semantic logic retrieval can mitigate the gap between potential visual logic in video frames and the logic expressed in the query. To be specific, constructing ternary logic triplets with visual elements (e.g., object1, logic type, object2) can enhance downstream reasoning capabilities when performing textual-visual retrieval.
@@
To the best of our knowledge, we are arguably the first to search for keyframes in long videos by detecting visual semantic logic, with potential extensions to other textual-visual retrieval tasks. Our main contributions are as follows:
@@
  • We define four fundamental types of semantic logic relations in video QA tasks, including temporal, causal, attribute, and spatial relations, which can be accurately detected across various datasets.
@@
  • We sample only 1.4% of frames on average of frames on average during keyframe search through heuristic sampling and distribution updating by different visual semantics and logical relations.
@@
  • We comprehensively evaluate retrieval efficiency, semantic similarity, temporal coverage, and video question answering accuracy across several widely used video understanding datasets, demonstrating significant improvements in downstream tasks.
@@ 2 Method
Figure 2: Our VSLS Framework for Efficient Keyframe Selection. VSLS sparsely samples frames and selects key ones via object detection and logic verification. Steps: 1) Use LLM&VLM to extract cue/target objects and four logic types (spatial, temporal, attribute, causal); 2) Adaptive sampling with evolving confidence; 3) Detect objects via YOLO-World; 4) Fuse scores with a spline function to identify high-confidence frames for downstream tasks.
Although existing long-context VLM frameworks implement keyframe search for video QA tasks 35; 47; 57; 65; 68; 80, their computational efficiency and searching accuracy remain suboptimal. To address this needle-in-a-haystack challenge 61; zhao2024needle, we propose a novel method VSLS that aligns the semantic relations between the text modality and video modality, enhancing the plausibility of logical reasoning and performance of downstream tasks.
@@ 2.1 Task Formulation
Given a video sequence $V=\{f_{t}\}_{t=1}^{N_{v}}$ with $N_{v}$ frames and a query $Q$, the ideal temporal search framework aims to retrieve the minimal keyframe subset $V^{K}=\{f_{m_{i}}\}_{i=1}^{K}\subseteq V$ with $K$ keyframes that satisfies:
@@
  • Compactness: $V^{K}$ must be a minimal subset that preserves completeness, which means that no frame in $V^{K}$ can be removed without hindering the accuracy and efficiency of video QA.
@@ 2.2 Visual Semantic Logic Extraction
Starting from a question $Q$ and uniformly sampled frames $\overline{V}_{N}$ from video $V$, our goal is to extract key visual elements to answer $Q$. We first classify the detected objects in $Q$ and $\overline{V}_{N}$ into two categories:
@@
To further leverage semantic and logical links among these objects, we define a set of relations $\mathcal{R}\subseteq\mathcal{O}\times\Delta\times\mathcal{O}$, where each relation $r=(o_{i},\delta,o_{j})\in\mathcal{R}$, with $o_{i},o_{j}\in\mathcal{O}$ denoting detected objects in the key and cue objects dataset, and $\delta\in\Delta$ representing one of the following types of relations:
[…]
$o_{i}$ and $o_{j}$ appear in the same frame, indicating co-occurrence or proximity.
Example: “A person is standing beside a vase.” $\Rightarrow$ (person, spatial, vase)
$o_{i}$ and $o_{j}$ share visual properties, e.g., color or size.
Example: “A person wears a black shirt.” $\Rightarrow$ (person, attribute, black shirt)
[…]
$o_{i}$ and $o_{j}$ occur in close frames, linking sequences or transitions.
Example: “After a dog entered the room, a cat entered.” $\Rightarrow$ (dog, temporal, cat)
$o_{i}$ and $o_{j}$ follow a cause-effect or prerequisite order.
Example: “A little girl broke the vase.” $\Rightarrow$ (little girl, causal, pieces)
The choice of these four relations draws on core concepts in linguistics and logic cohen1968universals; sowa2000knowledge; talmy2000toward, which identify spatial, temporal, attributive, and causal aspects as fundamental for structuring, perceiving, and communicating information about events and states. For more details on this selection, please see appendix LABEL:Theoretical for reference. As shown in Figure 1, we construct semantic-logical relations that support a broad range of question-answering tasks. Specifically, questions involving temporal queries (when does X happen?”), causal reasoning (why did Y occur?”), attribute dependence (What is the person wearing sunglasses doing?”), or spatial constraints (Who is standing next to the red car?”) can be answered more reliably by incorporating these structured relations and contextual cues.
@@ 2.3 Iterative Semantic-Logical Temporal Search
Based on the extracted key and cue objects and their logic relations, our algorithm iteratively searches for keyframes through semantic and logical reasoning, including four main stages: Frame Sampling (Sec. 2.3.1), Object Detection and Scoring (Sec. 2.3.2), Visual Semantic Logic Detection (Sec. 2.3.3), and Distribution Update (Sec. 2.3.4). The pseudocode is shown in Algorithm 1, and Algorithm LABEL:alg:slt_search provides a more detailed version.
Function SemanticLogicalTemporalSearch($V,Q,K,\Delta_{t},\tau,\alpha,\gamma$)
       $\mathcal{O},\mathcal{R}\leftarrow\text{ParseQuestion}(Q)$ // Extract key/cue objects and relations
      $P\leftarrow\text{Uniform},B\leftarrow|V|,S\leftarrow\emptyset,N_{v}\leftarrow|V|$ // Initialize distribution and state
      while $B>0$ and $|\mathcal{O}|>0$ do
             $k\leftarrow\lfloor\sqrt{B}\rfloor$, $G\leftarrow\textnormal{{Grid}}(\textnormal{{Sample}}(P,k^{2}))$ // Adaptive grid sampling
            $\Omega\leftarrow\textnormal{{DetectObjects}}(G)$ // Detect objects in sampled frames
            foreach $t\in G$ do
                   $C_{t}\leftarrow\textnormal{{CalculateBaseScore}}(\Omega_{t})$ // Base detection confidence
                  foreach  $r_{type}\in\mathcal{R}$  do
                         $\delta\leftarrow\textnormal{{Processrelation}}(r_{type},\Omega,\Delta_{t},\tau,\alpha,\gamma)$ // relations require distinct processing
                        $C_{t}\leftarrow C_{t}+\delta$
                  $\textnormal{{UpdateScores}}(S,t,C_{t})$ // Update global score registry
            $\textnormal{{DiffuseScores}}(S,w)$ // Temporal context propagation
            $P\leftarrow\textnormal{{NormalizeDistribution}}(S)$, $B\leftarrow B-k^{2}$ // Update sampling distribution
            foreach $g\in\textnormal{{TopK}}(S,K)$ do
                   if $\Omega[g]\cap\mathcal{O}\neq\emptyset$ then
                         $\mathcal{O}\leftarrow\mathcal{O}\setminus\Omega[g]$ // Remove identified key objects
      return $\textnormal{{TopK}}(S,K)$ // Return top-K keyframes
@@ 2.3.1 Frame Sampling
To accelerate the search process, we avoid exhaustively scanning all $N_{v}$ video frames and instead employ a distributed sampling strategy. Let $N_{v}$ denote the total number of frames in the video, and $P$ be a uniformly initialized sampling distribution over all frames. The sampling process is then defined as:
@@
where $\mathrm{Sample}(\cdot,N_{s})$ selects a subset of $N_{s}$ frames according to the distribution $P\odot N_{v}$. To further leverage the detecting ability of YOLO, we stack the sampled frames into a $k\times k$ grid, which imposes a constraint on the sample size $N_{s}$. Specifically, we require:
@@ 2.3.2 Object Detection and Scoring
In this stage, we construct the detection search space by taking the union of both key objects and cue objects. For each iteration, we detect objects on the $N_{s}$ sampled frames using a lightweight model like YOLO-World yolo for high efficiency and score the frames based on detection confidence. Specifically, let $\Omega_{t}$ be the set of detected objects in the frame at time $t$, $c_{o}$ the confidence of each detected object, and $w_{o}$ the corresponding weight. We define the frame score as:
$C_{t}=\max_{o\in\Omega_{t}}\bigl{(}c_{o}\cdot w_{o}\bigr{)}.$ (3)
@@ 2.3.3 Visual Semantic Logic Detection
Beyond individual object detection and frame-level scoring, we refine each frame’s confidence score by modeling higher-order object relations. Let $\mathcal{R}$ be the set of relations, where each $r\in\mathcal{R}$ involves a pair $(o_{1},o_{2})$ and is labeled by a type $r_{\text{type}}$. Denote $C_{t}$ as the confidence score at time $t$, with a global scaling factor $\alpha$ and a relation-specific weight $\gamma_{r_{\text{type}}}$ controlling each logic type’s impact. The refined confidence $C_{t}^{(r)}$ after applying relation $r$ is defined as:
@@
Attribute Relation. An attribute relation is satisfied when $o_{1}$ and $o_{2}$ share sufficient bounding-box overlap in the same frame. Let $\mathrm{overlap}$ be the ratio of their intersection area to the minimum of their individual bounding-box areas. If the overlap ratio exceeds a predefined threshold $\tau$ ($\tau=0.5$ in our experimental setting), we increase the frame confidence:
@@
Time Relation. A time relation checks whether two objects appear in temporally close frames. Suppose $t_{i}$ and $t_{j}$ ($t_{i}\leq t_{j}$) are sampled such that $|\,t_{j}-t_{i}|<\Delta_{t}$, where $\Delta_{t}$ is a threshold (e.g. 5 frames in our experimental setting), if $o_{1}$ occurs in frame $t_{i}$ and $o_{2}$ in frame $t_{j}$, then both frames’ confidences are updated:
@@
Through this scoring mechanism, frames with detected relations will have greater confidence and are more likely to be retrieved as keyframes for the given query and video. We have also conducted hyperparameter search experiments, and find that $\alpha=0.3$ (from 0.3, 0.5, 0.7, 1.0) and $\gamma_{r_{\text{type}}}=0.5$ achieve the best results across different datasets.
2.3.4 Distribution Update
After each iteration of frame sampling, we merge newly obtained frame confidences into the global score distribution $\{S_{f}\}$ spanning all frames $f=1,2,\ldots,N_{v}$. When a frame $f$ is selected for detection, its score is assigned to the confidence value $C_{f}$, and the visitation counter $N_{v,f}$ is reset to 0. To incorporate temporal context, we diffuse this updated score to neighboring frames within a window of size $w$. Denoting each nearby index by $f\pm\delta$ (for $\delta\in[-w,w]$), we apply: